### v2

# Infinite-Dimensional Closed-Loop Inverse Kinematics for Soft Robots via Neural Operators

Abstract
For fully actuated rigid robots, kinematic inversion is a purely geometric problem, efficiently solved by closed-loop inverse kinematics (CLIK) schemes that compute joint configurations to position the robot body in space. For underactuated soft robots, however, not all configurations are attainable through control action, making kinematic inversion extremely challenging. Extensions of CLIK address this by introducing end-to-end mappings from actuation to task space for the controller to operate on, but typically assume finite dimensions of the underlying virtual configuration space. In this work, we formulate CLIK in the infinite-dimensional domain to reason about the entire soft robot shape while solving tasks. We do this by composing an actuation-to-shape map with a shape-to-task map, deriving the differential end-to-end kinematics via an infinite-dimensional chain rule, and thereby obtaining a Jacobian-based CLIK algorithm. Since this actuation-to-shape mapping is rarely available in closed form, we propose to learn it using differentiable neural operator networks. We first present an analytical study on a constant-curvature segment, and then apply the neural version of the algorithm to a three-fiber soft robotic arm whose underlying model relies on morphoelasticity and active filament theory.
I Introduction
Kinematic inversion refers to finding a feasible robot configuration such that the robot assumes a desired positioning of one or more of its parts in space. For a rigid manipulator, this translates to computing joint positions that result in a desired end-effector pose, given the geometric link parameters. This formulation, however, assumes full actuation of rigid robots. When a robot is soft, it can bend and twist continuously rather than only at discrete joints. This flexibility enables navigation in complex environments and safe interaction with delicate objects, but it also makes control, even basic kinematic inversion, challenging.
For fully actuated systems, Jacobian-based closed-loop inverse kinematics (CLIK) schemes exploit the geometric freedom to select joint configurations directly, by using feedback in task-space to map back to required joint velocities via the Jacobian inverse [11, 16, 7]. However, in under-actuated systems with fewer degrees of actuation than degrees of freedom, such as soft robots, not all configurations are attainable through control action. Static kinematic constraints must be enforced explicitly, typically via a numerical optimization, which is then embedded, for example, in a model predictive control fashion [25, 24]. A compact, closed-loop form similar to the fully actuated case only becomes possible once we introduce an end-to-end kinematic mapping from the actuator inputs to the task-space output [26]: This mapping plays the role of a virtual joint-space model, like a controllable subspace, on which standard CLIK algorithms can then operate [5, 10, 21]. This so-called actuator-to-task closed-loop inverse kinematics approach was formalized for underactuated mechanical systems in [9]. Yet, still assuming finite dimensions for the underlying virtual joint space. i.e., relying on the common soft robot approximation that the shape is linearly parametrizable with a set of basis functions.
Figure 1: (Neural) Infinite-dimensional closed-loop inverse kinematics (CLIK): We propose an extension of finite-dimensional CLIK that allows us to reason on the whole soft robot shape while solving tasks. Since analytical expressions of the Jacobian are difficult to obtain in practice for infinite-dimensional soft robot models, we embed a neural version of the algorithm that uses the neural operator of the robot learned from simulations.
However, many soft-robot tasks require reasoning about the entire body shape, such as positioning any suitable point to a target or enforcing contact constraints along the body. In this work, we extend CLIK to systems with infinite-dimensional shape spaces, while keeping the controller finite-dimensional through end-to-end mappings from finite actuators to finite task dimensions. To this end, we define two mappings: (i) the mapping from actuators to shape, that is, the “internal” model of the (soft) robot, (ii) the forward kinematics from an infinite-dimensional shape space to a task, whose composition yields the end-to-end forward kinematics from actuation to task. Combining the gradient operators of both mappings via an infinite-dimensional chain rule yields the end-to-end differential forward kinematics, enabling the formulation of Jacobian-based CLIK. Since soft robots’ actuator-to-shape mappings often lack closed-form expressions or explicit Jacobians, this is only practical with a learned, differentiable model of this map. To address this, we propose to learn the mapping from simulation data using a neural operator network, which, unlike traditional neural networks, learns mappings between infinite-dimensional function spaces without fixed discretization, while providing well-defined gradients with respect to the finite-dimensional inputs [1, 18]. Neural operators are a recent trend in machine learning that have found wide interest in control and (soft) robotics already, for example as surrogate models for complex dynamics [4, 12, 19], or to make computationally expensive control schemes either real-time feasible or differentiable [2, 8, 13]. For us, they enable reasoning about the full infinite-dimensional soft robot shape during tasks.
II Preliminaries: Finite-Dimensional Closed-Loop Inverse Kinematics
We will briefly review the finite-dimensional case and unify notations, but refer to [22, 9] for details.
Kinematic inversion. Let $q\in\mathbb{R}^{n}$ be the configuration space of a mechanical system with $n$ degrees of freedom. Further, let $x\in\mathbb{R}^{p}$ be a $p$-dimensional task coordinate, commonly the end-effector position and orientation, such that the robot fulfilling the tasks is expressed through the forward kinematics
$\Phi_{\text{fin}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{p},\ q\mapsto x=\Phi_{\text{fin}}(q).$ (1)
The goal of kinematic inversion is to, given a desired task coordinate $\bar{x}$, find a configuration $\bar{q}$ such that $\bar{x}=\Phi_{\text{fin}}(\bar{q})$.
Closed-loop inverse kinematics (CLIK). CLIK solves the kinematic inversion in the form of dynamical system
$\dot{q}=c_{\text{IK}}(q,\bar{x})\qquad\text{s.t.}\lim_{t\to\infty}\Phi_{\text{fin}}(q(t))=\bar{x}.$ (2)
A standard choice that ensures exponential convergence under mild assumptions is
$\dot{q}=J_{\Phi_{\text{fin}}}(q)^{-1}K(\bar{x}-\Phi_{\text{fin}}(q))$ (3)
with positive definite gain matrix $K$ and Jacobian $J_{\Phi_{\text{fin}}}$, which is square and invertible for the fully actuated case.
CLIK for underactuated systems. For underactuated systems, we cannot control the $n$-dimensional configuration space, but we can extend the above to finding a control $c_{\text{IK}}$ that also respects static constraints, i.e.,
$\dot{q}$ $=c_{\text{IK}}(q,\bar{x})$
$\text{s.t.}\lim_{t\to\infty}\Phi_{\text{fin}}(q(t))=\bar{x}\text{ and }\lim_{t\to\infty}q(t)\in\mathcal{E},$ (4)
where $\mathcal{E}$ is the attainable equilibria set. To realize this, we introduce an end-to-end mapping from actuators to task output that creates a virtual configuration subspace controllable by actuators on which CLIK can operate. Therefore, consider the finite-dimensional mapping
$\Lambda_{\text{fin}}:\mathbb{R}^{m}\to\mathbb{R}^{n},\ \ q_{\text{a}}\mapsto q,$ (5)
from actuation coordinates $q_{\text{a}}$ with $m$ independent degrees of actuation to configuration coordinates $q$, such that the composed forward pseudo-kinematics are
$(\Phi\circ\Lambda)_{\text{fin}}:\ \mathbb{R}^{m}$ $\to\mathbb{R}^{p},$
$q_{\text{a}}$ $\mapsto x=(\Phi\circ\Lambda)_{\text{fin}}(q_{\text{a}}).$ (6)
Considering a square inversion ($m=p$) with Jacobian $J_{(\Phi\circ\Lambda)_{\text{fin}}}(q_{\text{a}})$, the standard gradient-based CLIK algorithm is
$\dot{q}_{\text{a}}=\left(J_{(\Phi\circ\Lambda)_{\text{fin}}}(q_{\text{a}})\right)^{-1}K(\bar{x}-(\Phi\circ\Lambda)_{\text{fin}}(q_{\text{a}})).$ (7)
A positive definite gain matrix $K\in\mathbb{R}^{m\times m}$ ensures exponential convergence analogous to the fully actuated case.
III Infinite-Dimensional Closed-Loop Inverse Kinematics
For the infinite-dimensional analogue of (7), we introduce a (soft) robot model via the actuation-to-shape mapping
$\Lambda:\mathbb{R}^{m}\to X,\ \ q_{\text{a}}\mapsto r=\Lambda(q_{\text{a}}),$ (8)
where $X=L^{2}([0,1];\mathbb{R}^{3})$ is a Hilbert space of square-integrable curves with respect to the normalized spatial coordinate $s\in[0,1]$, and equipped with the standard $L^{2}$-inner product. Furthermore, $r:[0,1]\to\mathbb{R}^{3}$ is a $C^{1}$ curve representing the soft robot shape in space [^1] Normalized arc-length $s\in[0,1]$ is used without loss of generality via the standard reparameterization $s=\tilde{s}/L$. The formulation also holds in $\mathbb{R}^{2}$.. Note that we parametrize attainable shapes by actuator coordinates $q_{\text{a}}\in\mathbb{R}^{m}$, so the attainable-equilibria manifold forms an $m$-dimensional submanifold of the infinite-dimensional function space $X$. Next, let $\Phi$ denote a $p$-dimensional task, i.e. the shape-to-task mapping
$\Phi:X\rightarrow\mathbb{R}^{p},\quad r\mapsto x=\Phi(r).$ (9)
such that the end-to-end mapping from actuation to task on which CLIK operates is the composition
$\Phi\circ\Lambda:\mathbb{R}^{m}\rightarrow\mathbb{R}^{p},\ \ q_{\text{a}}\mapsto x=(\Phi\circ\Lambda)(q_{\text{a}}).$ (10)
III-A Differential Kinematics
To realize the infinite-dimensional CLIK version of (7), we need the end-to-end Jacobian $J_{\Phi\circ\Lambda}(q_{\text{a}})\in\mathbb{R}^{p\times m}$, which is obtained via the differentials of the component mappings $\Lambda$ and $\Phi$, connected by the chain rule.
Actuation-to-shape. The Fréchet derivative of $\Lambda$ at $q_{\text{a}}$, $\Lambda^{\prime}(q_{\text{a}})$, is a bounded linear map
$\Lambda^{\prime}(q_{\text{a}}):\mathbb{R}^{m}\mapsto X,$
$\quad\Lambda(q_{\text{a}}+\delta q_{\text{a}})=\Lambda(q_{\text{a}})+\Lambda^{\prime}(q_{\text{a}})[\delta q_{\text{a}}]+o(\|\delta q_{\text{a}}\|),$ (11)
that maps small changes in actuators to small changes in shape. Since our actuation is finite dimensional, the Fréchet derivative reduces to a linear combination of partial derivatives in the function space,
$\Lambda^{\prime}(q_{\text{a}})[\delta q_{\text{a}}]=\sum_{i=1}^{m}\delta q_{\text{a},i}\frac{\partial\Lambda(q_{\text{a}})}{\partial q_{\text{a},i}},$ (12)
$\quad\frac{\partial\Lambda(q_{\text{a}})}{\partial q_{\text{a},i}}=:\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})\in X.$
Figure 2: Block diagram of the infinite-dimensional closed-loop inverse kinematics (23) in actuation space $q_{\text{a}}\in\mathbb{R}^{m}$, with a square actuation-to-task mapping. Trapezoids emphasize the change in dimensionality from input to output.
Shape-to-task. For a perturbation $v:[0,L]\mapsto\mathbb{R}^{3}$, define the perturbed shape $r_{\varepsilon}=r+\varepsilon v$ and the directional derivative
$\Phi^{\prime}(r)[v]=\frac{d}{d\varepsilon}\Phi(r_{\varepsilon})\big|_{\varepsilon=0}.$ (13)
The directional derivative $\Phi^{\prime}(r)[v]$ defines a bounded linear functional on the tangent space of admissible perturbations $v$. Endowing this space with a Hilbert structure (such as $X=L^{2}([0,1];\mathbb{R}^{3})$) with inner product
$\langle u,v\rangle_{L^{2}}=\int_{0}^{1}u(z)\cdot v(z)dz,$ (14)
the Riesz representation theorem [20] ensures the existence of a unique element $\nabla_{X}\Phi(r)\in X$ such that
$\Phi^{\prime}(r)[v]=\langle\nabla_{X}\Phi(r),v\rangle_{X},\quad\forall v\in X,$ (15)
i.e. the $L^{2}$ gradient of $\Phi$ at $r$.
End-to-end. For the CLIK algorithm, we consider the differential end-to-end kinematics
$\dot{x}=J_{\Phi\circ\Lambda}(q_{\text{a}})\dot{q}_{\text{a}},\ J_{\Phi\circ\Lambda}\in\mathbb{R}^{p\times m}.$ (16)
For any direction $\dot{q}_{\text{a}}\in\mathbb{R}^{m}$, by chain rule, the derivative of this composition is
$(\Phi\circ\Lambda)^{\prime}(q_{\text{a}})[\dot{q}_{\text{a}}]=\big(\Phi^{\prime}(r)\big)\big(\Lambda^{\prime}(q_{\text{a}})[\dot{q}_{\text{a}}]\big).$ (17)
Together with the task differential $\Phi^{\prime}(r)[v]$ (15), this yields
$(\Phi\circ\Lambda)^{\prime}(q_{\text{a}})[\dot{q}_{\text{a}}]$ $=\big\langle\nabla_{X}\Phi(r),\;\Lambda^{\prime}(q_{\text{a}})[\dot{q}_{\text{a}}]\big\rangle_{X}$ (18)
$=\big\langle\Lambda^{\prime}(q_{\text{a}})^{*}\,\nabla_{X}\Phi(r),\;\dot{q}_{\text{a}}\big\rangle_{\mathbb{R}^{m}},$ (19)
where $\Lambda^{\prime}(q_{\text{a}})^{*}:X\rightarrow\mathbb{R}^{m}$ is the Hilbert adjoint. Comparing this with the differential end-to-end kinematics (16), we note that the Jacobian of the composition must be
$J_{\Phi\circ\Lambda}(q_{\text{a}})=\Lambda^{\prime}(q_{\text{a}})^{*}\nabla_{X}\Phi(\Lambda(q_{\text{a}}))\in\mathbb{R}^{p\times m},$ (20)
and its $i$-th column is calculated as
$J_{\Phi\circ\Lambda,:,i}(q_{\text{a}})$ $=\langle\nabla_{X}\Phi(\Lambda(q_{\text{a}})),\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})\rangle_{X},$ (21)
$=\int_{0}^{L}\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})(z)\cdot\nabla_{X}\Phi(\Lambda(q_{\text{a}}))(z)\ dz$ (22)
for $i=1,2,..,m$. When $p=m$ and $J_{\Phi\circ\Lambda}(q_{\text{a}})$ is nonsingular, following the usual steps for kinematic inversion yields the final CLIK algorithm (cf. Figure 2)
$\dot{q}_{\text{a}}$ $=\left(J_{\Phi\circ\Lambda}(q_{\text{a}})\right)^{-1}K(\bar{x}-(\Phi\circ\Lambda)(q_{\text{a}})),$ (23a)
$r$ $=\bar{r}(q_{\text{a}})$ (23b)
with positive definite gain matrix $K\in\mathbb{R}^{m\times m}$.
Proposition 1
The closed-loop kinematic inversion (23) solves (4) exponentially fast $\forall K\succ 0$.
Proof:
First, we formulate the differential kinematics of the end-to-end mapping (16), and plug in the proposed controller (23)
$\dot{x}$ $=J_{\Phi\circ\Lambda}(q_{\text{a}})\dot{q}_{\text{a}}$ (24)
$=J_{\Phi\circ\Lambda}(q_{\text{a}})\left(J_{\Phi\circ\Lambda}(q_{\text{a}})\right)^{-1}K(\bar{x}-(\Phi\circ\Lambda)(q_{\text{a}}))$ (25)
$=K(\bar{x}-(\Phi\circ\Lambda)(q_{\text{a}})).$ (26)
We observe that the resulting error dynamics $\dot{\bar{x}}-\dot{x}=-K(\bar{x}-x)$ is linear and converges exponentially fast to zero with the rate defined by $K$. Consequently, the steady state of (23) fulfills $\lim_{t\to\infty}\Phi(q(t))=\bar{x}$. ∎
III-B Representative Tasks and Resulting Jacobians
We can think of many functionals $\Phi:X\to\mathbb{R}^{p}$ in the task space of a robot. The most common examples involve reaching a target or a prescribed distance from it for a specific point along the robot’s body, typically the end-effector for rigid robots. While such tasks are equally relevant for soft robots, their continuous nature invites an adaptation of this formulation: rather than requiring the end-effector to reach the target, we may instead define the task such that the point along the robot’s body closest to the target reaches it (cf. Figure 3).
Positioning tasks. Reaching a target point $x_{0}$ (or time-varying $x_{0}(t)$) at a fixed centerline coordinate $\bar{s}\in[0,1]$ (or $\bar{s}(t)$) of the robot centerline $r$ translates to
$\Phi_{\text{pos}}^{\text{fixed}}(r)$ $=r(\bar{s})-x_{0}.$ (27)
However, our infinite-dimensional CLIK shines for the more expressive closest-point task, which automatically identifies the best location along the entire robot shape to reach the target. This is formulated as
$\Phi_{\text{pos}}^{\text{opt}}(r)$ $=r(s_{*})-x_{0},$ (28a)
$\text{where }s_{*}$ $=\text{argmin}_{s\in[0,1]}\frac{1}{2}\|r(s)-x_{0}\|^{2},$ (28b)
meaning that the closest centerline coordinate $s_{*}$ to the target is computed via gradient descent in each step.
The $L^{2}$-gradients of the positioning tasks are $\nabla_{L^{2}}\Phi_{\text{pos}}^{\text{fixed}}(r)=\delta(s-\bar{s})\in X$, and $\nabla_{L^{2}}\Phi_{\text{pos}}^{\text{opt}}(r)=\delta(s-s_{*})\in X$, such that the Jacobians of the end-to-end mapping $J_{\Phi\circ\Lambda}\in\mathbb{R}^{3\times m}$ have $m$ columns
$J_{(\Phi_{\text{pos}}^{\text{fixed}}\circ\Lambda):,i}(q_{\text{a}})=\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})(\bar{s}),$ (29)
$J_{(\Phi_{\text{pos}}^{\text{opt}}\circ\Lambda):,i}(q_{\text{a}})=\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})(s_{*}),$ (30)
i.e., only the partial derivative of the actuation-to-shape mapping at the respective $s$-coordinate remains.
Distance tasks. Scalar distance tasks follow the same structure. The fixed-coordinate variant minimizes squared distance from a prescribed centerline coordinate $\bar{s}\in[0,1]$
$\Phi_{\text{dist}}^{\text{fixed}}(r)$ $=\frac{1}{2}\|r(\bar{s})-x_{0}\|^{2}.$ (31)
Again, infinite-dimensional CLIK excels when minimizing the distance from the closest point on the centerline
$\Phi_{\text{dist}}^{\text{opt}}(r)$ $=\min_{s\in[0,1]}\frac{1}{2}\|r(s)-x_{0}\|^{2},$ (32)
where, the closest centerline coordinate $s_{*}$ is computed via gradient descent. Assuming uniqueness and nondegeneracy of the minimizer (Infinite-dimensional Danskin’s Theorem, cf. Theorem 4.13 in [3]), its directional derivative follows (15) with $L^{2}$-gradients $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{fixed}}(r)=(r(s_{*})-x_{0})\delta(s-s_{*})\in X$, and $\nabla_{L^{2}}\Phi_{\text{dist}}^{\text{opt}}(r)=(r(s_{*})-x_{0})\delta(s-\bar{s})\in X$, such that the row Jacobians of the end-to-end mapping $J_{\Phi\circ\Lambda}\in\mathbb{R}^{1\times m}$ have $m$ scalar entries
$J_{(\Phi_{\text{dist}}^{\text{fixed}}\circ\Lambda),i}(q_{\text{a}})$ $=\left[\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})(\bar{s})\right]^{\top}(r(\bar{s})-x_{0}),$ (33)
$J_{(\Phi_{\text{dist}}^{\text{opt}}\circ\Lambda),i}(q_{\text{a}})$ $=\left[\partial_{q_{\text{a},i}}\Lambda(q_{\text{a}})(s_{*})\right]^{\top}(r(s_{*})-x_{0}).$ (34)
Figure 3: While standard tasks usually focus to position the end-effector $r(\bar{s}=1)$, the continuum nature of the soft robot invites for just reaching a target point with the closest point $r(s_{*})$ on its backbone, where $s_{*}$ is adapted in each iteration.
III-C Toy Example: Constant Curvature Segment
Let us now consider an analytical example to demonstrate the infinite-dimensional CLIK algorithm. For the sake of simplicity, and almost w.l.o.g., assume a planar, inextensible robot such that the only strain is the curvature $\kappa(s)$, where $s\in[0,1]$ is a normalized arc‑length coordinate and the physical arc length is $L$. This leads to infinite-dimensional mapping $\Lambda_{\kappa}$ that maps curvature functions to a backbone pose ($x,y,\alpha$), more precisely,
$\Lambda_{\kappa}(\kappa)(s)=\begin{bmatrix}x(s)\\ y(s)\\ \alpha(s)\end{bmatrix}=\begin{bmatrix}L\int_{0}^{s}\cos(\alpha(v))dv\\ L\int_{0}^{s}\sin(\alpha(v))dv\\ \int_{0}^{s}\kappa(v)dv\end{bmatrix}$. (35)
Assuming a constant curvature $q$ along the segment such that $\kappa(s)\approx q$, and focusing on the centerline position $r(s)=[x(s),y(s)]$ without orientation, this simplifies to
$\Lambda_{\text{CC}}(q)(s)=r(s)=\begin{bmatrix}L\frac{\sin(sq)}{q}\\ L\frac{1-\cos(sq)}{q}\end{bmatrix}$ (36)
with Jacobian
$J_{\text{CC}}(q)=L\begin{bmatrix}\frac{sq\cos(sq)-\sin(sq)}{q^{2}}&\frac{(\cos(sq)-1)+sq\sin(sq)}{q^{2}}\end{bmatrix}^{\top}$. (37)
Since the constant curvature segment has only one degree of actuation, namely this curvature itself, i.e. $q_{\text{a}}=q$, we require a one dimensional task such as the scalar distance task $\Phi_{\text{dist}}^{\text{fixed}}$. The scalar Jacobian of the composition is then
$J_{\Phi_{\text{dist}}^{\text{fixed}}\circ\Lambda_{\text{CC}}}(q)=J_{\text{CC}}(q)(\bar{s})(r(\bar{s})-x_{0}).$ (38)
Reaching a point translates to eliminating the distance, i.e. $\bar{x}=\Phi_{\text{dist}}^{\text{fixed}}(\bar{r})=0$, such that the CLIK control is
$\dot{q}=-K\left[J_{(\Phi_{\text{dist}}^{\text{fixed}}\circ\Lambda_{\text{CC}})}(q)\right]^{-1}\Phi_{\text{dist}}^{\text{fixed}}(\Lambda_{\text{CC}}(q))$ (39)
with scalar weight $K>0$. The same holds for the closest-point version of the task $\Phi_{\text{dist}}^{\text{opt}}$ with $s_{*}$ instead of $\bar{s}$.
Figure 4 shows simulations of such a constant curvature segment reaching a target point at a fixed backbone coordinate (the tip position in this case), as well as the closest-point task where any point on the robot should reach the target as fast as possible. In both scenarios, the scalar weight is $K=10$ and we observe exponential convergence. For the closest-point task, the optimal coordinate $s_{*}$ is continuously updated to the point nearest to the target.
Figure 4: Constant curvature segment: Robot evolution and time evolutions ($t\in[0,1]$) of the actuation variable $q$ and the task variable $x$ for the point-reaching task $\Phi_{\text{pos}}^{\text{fixed}}$ (top) and the closest-point task $\Phi_{\text{opt}}^{\text{fixed}}$ (bottom).
IV Neural Closed-Loop Inverse Kinematics
While the previous example demonstrates how the infinite-dimensional CLIK works and what possibilities we have when being able to reason on the whole soft body shape while solving tasks, we recognize that the first of the two maps $\Lambda$ is very hard to obtain in closed form in practice, especially once more complex soft robots are involved. This is why we propose to learn the soft robot model operator of the actuation-to-shape mapping $\Lambda$ (8), denoted as $\mathcal{G}_{\Lambda}$, via neural operator networks.
IV-A Neural Operator Networks and Their Gradients
The universal operator theorem [6] shows that a neural network with one hidden layer can approximate any continuous nonlinear operators. Extending this, the DeepONet [18] architecture allows for learning operators of the form $\mathcal{G}:u\mapsto\mathcal{G}(u)$, where both $u$ and $\mathcal{G}(u)$ are functions defined on (infinite-dimensional) spaces. For any point $s$ in the domain of $\mathcal{G}(u)$, the value $\mathcal{G}(u)(s)$ is a real-valued vector in $\mathbb{R}^{d}$. To learn said operator, DeepONet takes as inputs the sampled values $u(x_{1}),u(x_{2}),...,u(x_{m})]$ at a finite number of fixed locations $x_{1},x_{2},...,x_{m}$}, together with the evaluation points $s$. The architecture consists of two sub-networks, the branch network processing the input function, and the trunk network processing the evaluations points,
$\mathcal{G}^{\text{branch}}$ $:\mathbb{R}^{m}\to\mathbb{R}^{v\times d},\quad u\mapsto a_{\text{br}},$ (40)
$\mathcal{G}^{\text{trunk}}$ $:\mathbb{R}\to\mathbb{R}^{v\times d},\quad s\mapsto a_{\text{tr}},$
where $v$ is a hyperparameter controlling the width of the latent representation. The final prediction is computed as the inner product of the outputs of the branch and trunk networks over the latent dimension
$\mathcal{G}(u)(s)$ $=a_{\text{br}}\odot a_{\text{tr}},$ (41)
where $\odot$ denotes the element-wise multiplication over the latent dimension. Once trained, the DeepONet realization is a differentiable mapping with respect to all its inputs, so gradients with respect to both the input function samples and the evaluation coordinate $s$ are obtained directly by automatic differentiation.
In practice, the branch network is trained on finitely many function samples, while at inference time the trunk network can be evaluated at arbitrarily many, arbitrarily fine locations $s$. This decouples the discretization of the input function from the resolution of the output, which is a key advantage of neural operators over discretization-based neural networks, and makes it particularly powerful for infinite-dimensional CLIK where we want to reason over the continuous robot shape and might be interested in finer discretizations.
IV-B Example: Three-fiber Soft Robotic Arm
To demonstrate our neural infinite-dimensional CLIK algorithm, we consider a soft robotic arm initially designed with three contractable fibers, inspired by the elephant trunk musculature. It is depicted in Figure 5. By design, this structure is soft and slender, and can grow, shrink, or actively remodel its intrinsic shape while also deforming elastically – which is why it is modelled as an active, morphoelastic filament [15]. This model has been previously validated in experiments [17, 14], successfully used for path-planning algorithms [23], and shall now serve as example for the neural operator implementation of infinite-dimensional CLIK. Its forward kinematics are governed by
$r^{\prime}(s)$ $=\zeta d_{3},$ (42a)
$d_{i}^{\prime}(s)$ $=\hat{\zeta}u\times d_{i},\quad i=\{1,2,3\},$ (42b)
$0$ $=\frac{\partial n}{\partial s}+\hat{\zeta}f$ (42c)
$0$ $=\frac{\partial m}{\partial s}+\frac{\partial r}{\partial s}\times n+\hat{\zeta}l.$ (42d)
with centerline $r:[0,1]\rightarrow\mathbb{R}^{3}$, normalized spatial coordinate $s\in[0,1]$, orthonormal director frame $D=\{d_{1},d_{2},d_{3}\}$, as well as extensions $\zeta$, curvatures $u$, internal force $n$, external force $f$, internal momentum $m$. The mapping from actuation $q_{\text{a}}\in\mathbb{R}^{3}$ to shape $r(s)$ is mediated by the intrinsic curvature $\hat{u}$ and extension $\hat{\zeta}$, both of which are functions of the fiber activations, hence, the actuation, i.e., $\hat{u}(q_{\text{a}})$ and $\hat{\zeta}(q_{\text{a}})$. Details are found in [15, 14].
IV-C Learning the Model with a Neural Operator
In our application, the architecture of the operator network is simplified compared to DeepONet [18], since the actuation coordinates $q_{\text{a}}\in\mathbb{R}^{m}$ are finite-dimensional. Consequently, the branch net reduces to a standard feedforward network, while the trunk net only encodes the spatial variable $s$. We still retain the key property that the operator is learned over actuation and space, so we can evaluate the learned actuation-to-shape mapping $\mathcal{G}_{\Lambda}$ at arbitrarily many points $s$ and differentiate with respect to $q_{\text{a}}$.
Figure 5: Three-fiber soft robotic arm [14, 17]: The orientation of the fibers is inspired by the muscular structure of the elephant trunk, with one straight fiber and a helical fiber pair, enabling continuous deformations and optimal reachability.
Data. For data generation, we solve the boundary value problem (42) $N=1,000,000$ times to obtain samples of the actuation-to-shape mapping for random activations within the physical constraints of the system $q_{\text{a},i}\in[-1.67,0]$, $i=1,2,3$. Note that $q_{\text{a}}=0$ corresponds to the steady state of the system in which the soft robot hangs downward, and negative actuation leads to fiber contraction relative to this state. Histograms of the activations show that the entire actuation space is sufficiently represented in the dataset. The spatial coordinate $s$ is discretized with $n_{\text{s}}=100$ points in $[0,1]$. We use the Julia package developed for the active filament theory [15] and the associated parameters of the minimal three fiber prototype [17]. We store each solution as a discrete centerline in a matrix $R\in\mathbb{R}^{n_{s}\times 3}$ with its respective $(x,y,z)$-coordinates at the $n_{s}$ locations.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 6: Three-fiber soft robotic arm. Robot evolution (   ) and time evolutions of the actuation variable (activations $\gamma_{i}$) and the task variable $x$ for the fixed point-reaching task $\Phi_{\text{pos}}^{\text{fixed}}$ (left) closest-point task $\Phi_{\text{pos}}^{\text{opt}}$ (right). For the robot evolution, we plot every 10th shape.
Training process. Of the 1,000,000 samples, 200,000 are held out as a test set. From the remaining data, 640,000 samples are used for training and 160,000 for validation. We train the operator network for 500 epochs with the Adam optimizer, tanh activations, batch size 32, and an exponentially decaying learning rate, using the validation loss to select the best model checkpoint (typically reached after 400-450 epochs). On the held-out test set, the trained model achieves a mean squared error of $1.38\times 10^{-10}$ and an $L^{2}$ relative error of $6.08\times 10^{-4}$. The architecture and test performance are summarized in Table I.
Architecture and Training
Optimizer Adam
Activation tanh
Learning rate Exponential decay
Epochs 500
Batch size 32
Branch layers [3, 64, 64, 64, 192]
Trunk layers [1, 64, 64, 64, 192]
Test Performance
MSE $1.38\times 10^{-10}$
L2 relative error $6.08\times 10^{-4}$
TABLE I: Neural operator hyperparameters and performance.
IV-D Composed Jacobian and Neural CLIK
The soft robot has three actuators, hence, we consider the three-dimensional tasks $\Phi_{\text{pos}}^{\text{fixed}}$ and $\Phi_{\text{pos}}^{\text{opt}}$ that position the robot with respect to a target point in $(x,y,z)$ coordinates. From automatic differentiation, we obtain the gradient of the operator network $\mathcal{G}_{\Lambda}$ with respect to the input $q_{\text{a}}$. With the task gradients being Dirac at $\bar{s}$ or $s_{*}$, the Jacobian of the composition is the evaluation of the operator network gradient at the required spatial coordinate $\bar{s}$ or $s_{*}$ according to (29), (30), namely
$J_{\Phi_{\text{pos}}^{\text{fixed}}\circ\mathcal{G}_{\Lambda}}(q_{\text{a}})$ $=\frac{\partial\mathcal{G}_{\Lambda}(q_{\text{a}})}{\partial q_{\text{a}}}(\bar{s})\in\mathbb{R}^{3\times 3},$ (43)
$J_{\Phi_{\text{pos}}^{\text{opt}}\circ\mathcal{G}_{\Lambda}}(q_{\text{a}})$ $=\frac{\partial\mathcal{G}_{\Lambda}(q_{\text{a}})}{\partial q_{\text{a}}}(s_{*})\in\mathbb{R}^{3\times 3}.$ (44)
Reaching a point translates to a desired task variable $\bar{x}=\Phi_{\text{pos}}(\bar{r})=[0,0,0]$, and the control law follows as (23).
Figure 6 shows simulations of the neural version of the infinite-dimensional CLIK algorithm for the two distinct positioning tasks. The simulation is for $t\in[0,1]$ with a time step of $dt=0.001$. The weighting matrix is $K=8\cdot\text{diag}(1,1,1)$, i.e., all positioning coordinates are equally weighted, and the magnitude is a balance between convergence speed and overshooting. We use a spatial discretization with $n_{\text{s}}=100$ points in this example, but this discretization could be arbitrarily finer. The code and dataset will be made available on GitHub upon acceptance.
V Conclusion
In this paper we extend classical closed-loop inverse kinematics (CLIK) to the infinite-dimensional shape space of soft robots, enabling controllers to reason directly on the entire robot shape while solving tasks. By composing an actuation-to-shape map with a shape-to-task map and applying an infinite-dimensional chain rule, we derive the corresponding differential end-to-end kinematics and obtain a Jacobian-based CLIK algorithm. Since the actuation-to-shape mapping typically lacks closed-form expressions, we show how to learn it from simulation data using differentiable neural operator networks. The simulations focus on simple positioning tasks, yet the framework naturally extends to more complex objective functionals, and future work will explore how to incorporate obstacles or task-space trajectory tracking. Experimental validation is necessary to assess how the algorithm performs under model mismatch between the simulated (learned) operator and the real robot.